OpenStudy (anonymous):

Simplify: (3x2 − 2) + (5x2 + 5x − 1)

OpenStudy (anonymous):

for this you want to combine "like Terms"

OpenStudy (anonymous):

can you right it out and i do the work?

OpenStudy (anonymous):

sure you have a 3x^2 and a 5x^2, since they both have the similar term x^2 you can combine those you have a 5x, but nothing else with just an 'x' as a variable so that stays 5x and finally you have -2 + -1, so just add those terms

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so your final answer is?

OpenStudy (anonymous):

2x^2 +5x=-3

OpenStudy (anonymous):

i think the parenthesis are messing you up...what if you looked at it like this? 3x^2 + 5x^2= 5x -2 + -1=

OpenStudy (anonymous):

8x2 + 5x − 3? or 2x2 + 5x − 3?

OpenStudy (anonymous):

youre adding 3x^2 PLUS 5x^2 so 3+5= 8x^2 so youre final answer is: 8x^2 +5x -3

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

no problem, just remember when adding/subtracting to combine like terms!